OpenStudy (anonymous):

solve the equation y/5=-3/10

OpenStudy (anonymous):

y = −3/2

OpenStudy (anonymous):

what i'm asking is how do I find Y? I don't want to use my calculator

OpenStudy (anonymous):

y/5=-3/10 or, y= (-3/10)*5 = -15 / 10 = -3/2 r u getting the answer?

OpenStudy (anonymous):

i just said that -_-